OpenStudy (anonymous):

HELP ME!!!!!! What is the length of the conjugate axis? ((y-2)^2/16)-((x+1)^2/144)=1

OpenStudy (tkhunny):

24 - Eyeball problem. Don't let this kind get away from you.

OpenStudy (anonymous):

how did you get that?

OpenStudy (tkhunny):

There's a positive and a negative piece in this form. The positive piece (y in this case), listed first, defines the transverse axis in it's denominator. 16 = 4^2 and then 2*4 = 8 -- This is the length of the transverse axis Also, usually stated simply as a = 4 The negative piece (x in this case), listed second, defines the conjugate axis in it's denominator. 144 = 12^2 and then 2*12 = 24 -- This is the length of the conjugate axis Also, usually stated simply as b = 12

OpenStudy (anonymous):

ok i got it now thank you